# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ Appendix D Proof of Theorem 1
D.1 Main Proof of Theorem 1
@@
$\ell_{t}(\bm{\theta}):=\sum_{j=1}^{|S_{t}|}\underbrace{-\log\left(\frac{\exp\left(\phi(x_{t},\sigma_{tj})^{\top}\bm{\theta}\right)}{\sum_{k=j}^{|S_{t}|}\exp\left(\phi(x_{t},\sigma_{tk})^{\top}\bm{\theta})\right)}\right)}_{=:\ell^{(j)}_{t}(\bm{\theta})}=\sum_{j=1}^{|S_{t}|}\ell^{(j)}_{t}(\bm{\theta}).$
@@
Online confidence bound for PL loss. Now, we present the confidence bound for online parameter estimation in MNL models, as recently proposed by Lee and Oh [43].
Lemma D.1 (Online confidence bound, Theorem 4.2 of Lee and Oh 43).
@@
Corollary D.1 (Restatement of Corollary 1, Online confidence bound for PL loss).
@@
Useful definitions. We define the set of warm-up rounds, denoted by ${\mathcal{T}}^{w}$, which consists of rounds with large uncertainty, as follows:
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\right\},$ (D.1)
where $\beta_{T+1}(\delta)$ denotes the confidence radius as defined in Corollary D.1. Furthermore, we define the regularized sample covariance matrix of feature differences (with respect to $\phi(x_{s},\bar{a}_{s})$) over the non-warm-up rounds as:
$\Lambda_{t}:=\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}.$ (D.2)
To control the elliptical potentials, we also define the set of large elliptical potential (EP) rounds, denoted by ${\mathcal{T}}_{0}$, as follows:
${\mathcal{T}}_{0}$ $:=\left\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\right\},$ (D.3)
@@
Lemma D.2 (Restatement of Lemma 1).
Let $\Lambda_{t}$ be defined as in Equation (D.2). Set $\lambda=\Omega(d\log(KT/\delta))$. Then, for all $t\in[T]$, with probability at least $1-\delta$, we have
$H_{t}\succeq\frac{1}{50}\Lambda_{t}.$
The proof is deferred to Appendix D.2.1.
The following lemma is a variant of the elliptical potential lemma [1], adapted specifically to the assortment offering setting. For completeness, we provide the proof tailored to our setting.
Lemma D.3 (Elliptical potential lemma for $S_{t}$).
Let $\{z_{ta}\}_{t\geq 1,a\in S_{t}}$ be a bounded sequence in $\mathbb{R}^{d}$ satisfying $\max_{t\geq 1}\|z_{ta}\|_{2}\leq X$. For any $t\geq 1$, we define $\Lambda_{t}:=\sum_{s=1}^{t-1}\sum_{a\in S_{s}}z_{sa}z_{sa}^{\top}+\lambda\mathbf{I}_{d}$ with $\lambda>0$. Then, we have
$\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\right\}\leq 2d\log\left(1+\frac{X^{2}KT}{d\lambda}\right).$
The proof is deferred to Appendix D.2.2.
The cardinality of the set ${\mathcal{T}}_{0}$ can be bounded by a variant of the elliptical potential counting lemma [40, 35].
Lemma D.4 (Elliptical potential count lemma for $S_{t}$).
Let $\{z_{ta}\}_{t\geq 1,a\in S_{t}}$ be a bounded sequence in $\mathbb{R}^{d}$ satisfying $\max_{t\geq 1}\|z_{ta}\|_{2}\leq X$. For any $t\geq 1$, we define $\Lambda_{t}:=\sum_{s=1}^{t-1}\sum_{a\in S_{s}}z_{sa}z_{sa}^{\top}+\lambda\mathbf{I}_{d}$ with $\lambda>0$. Let ${\mathcal{T}}_{0}\subseteq[T]$ be the set of indices where $\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\geq L$. Then,
$|{\mathcal{T}}_{0}|\leq\frac{2d}{\log(1+L)}\log\left(1+\frac{X^{2}K}{\log(1+L)\lambda}\right).$
The proof is deferred to Appendix D.2.3.
The size of the set ${\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$ is bounded as described in the following lemma:
Lemma D.5.
Let ${\mathcal{T}}_{0}:=\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\}$ and ${\mathcal{T}}^{w}=\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\}$. Define $\kappa:=e^{-4B}$. Then, the size of the set ${\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$ is bounded as follows:
$\left|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}\right|\leq\frac{12\sqrt{2}K^{2}}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right).$
The proof is deferred to Appendix D.2.4.
@@
$\leq\frac{1}{T}\sum_{t=1}^{T}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right),$ (D.4)
@@
To complete the proof, it remains to bound the first term in Equation (D.4).
Recall the definitions of the set of large elliptical potential (EP) rounds (Equation (D.3)), denoted by ${\mathcal{T}}_{0}$, and the set of warm-up rounds (Equation (D.1)), denoted by ${\mathcal{T}}^{w}$:
${\mathcal{T}}_{0}$ $:=\left\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\right\},$ (large EP rounds)
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\right\},$ (warm-up rounds)
where $\Lambda_{t}$ is defined in Equation (D.2). Then, by applying the elliptical potential count lemma (Lemma D.4) and the bound on the cardinality of the set $|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}|$ lemma (Lemma D.5), we obtain
@@
$=\frac{1}{T}\sum_{t\in{\mathcal{T}}_{0}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
$+\frac{1}{T}\sum_{t\in{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{4B}{T}|{\mathcal{T}}_{0}|+\frac{4B}{T}|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}|+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$ (Assumption 1)
@@
$\leq\frac{8B}{\log(2)T}d\log\left(1+\frac{2K}{\log(2)\lambda}\right)+\frac{48\sqrt{2}BK^{2}}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right)$ (Lemma D.4 and D.5)
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right).$ (D.5)
To further bound the last term of Equation (D.5), we get
@@
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}$ $\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{T+1}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ (Hölder’s ineq.)
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ ($H_{T+1}\succeq H_{t}$)
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}.$ (Corollary D.1, with prob. $1-\delta$)
We denote $S^{\star}_{t}=\{\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})\}$. Then, we have
@@
$\frac{\beta_{T+1}(\delta)}{T}$ $\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$=\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$=\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{|S^{\star}_{t}|}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$=\frac{2\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}},$ (D.6)
where the last equality holds due to the fact that $|S^{\star}_{t}|=2$. To proceed, by our efficient assortment selection rule in Equation (9), we obtain
$\frac{2\beta_{T+1}(\delta)}{T}$ $\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}}$ ($S_{t}$ selection rule, Eqn. (9))

### v2

@@ Appendix C Proof of Theorem 1
C.1 Main Proof of Theorem 1
@@
$\ell_{t}(\bm{\theta}):=\sum_{j=1}^{|S_{t}|}\underbrace{-\log\left(\frac{\exp\left(\phi(x_{t},\sigma_{tj})^{\top}\bm{\theta}\right)}{\sum_{k=j}^{|S_{t}|}\exp\left(\phi(x_{t},\sigma_{tk})^{\top}\bm{\theta})\right)}\right)}_{=:\ell^{(j)}_{t}(\bm{\theta})}=\sum_{j=1}^{|S_{t}|}\ell^{(j)}_{t}(\bm{\theta}),$
@@
Online confidence bound for PL loss. Now, we present the confidence bound for online parameter estimation in MNL models, as recently proposed by lee2025improved.
Lemma C.1 (Online confidence bound, Theorem 4.2 of lee2025improved).
@@
Corollary C.1 (Restatement of Corollary 1, Online confidence bound for PL loss).
@@
Useful definitions. For clarity, we write $\ell_{t,\sigma}(\bm{\theta})$ to make explicit that the loss function is random with respect to the ranking $\sigma$. When the ranking is realized as $\sigma_{t}$, we simply denote it by $\ell_{t}(\bm{\theta})$. We define the set of warm-up rounds, denoted by ${\mathcal{T}}^{w}$, which consists of rounds with large uncertainty, as:
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}K\beta_{T+1}(\delta)}\right\},$ (C.1)
where $\beta_{T+1}(\delta)$ denotes the confidence radius as defined in Corollary C.1.
Furthermore, we define the expected version of the Hessian matrix (taken with respect to the randomness of the ranking feedback) as follows:
$M_{t}=\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})}\left[\nabla^{2}\ell^{(j)}_{s,\sigma}(\widehat{\bm{\theta}}_{s})\right]+\lambda\mathbf{I}_{d}$ (C.2)
Note that $\nabla^{2}\ell^{(j)}_{s,\sigma}(\widehat{\bm{\theta}}_{s})$ is a random matrix conditional on the assortment $S_{s}$, where the randomness arises from the ranking feedback $\sigma$.
@@
Lemma C.2 (Empirical-to-expected Hessian lower bound).
Let $M_{t}$ be defined as in Equation (C.2). Set $\lambda=\Omega(d\log(KT/\delta))$. Then, for all $t\in[T]$, with probability at least $1-\delta$, we have
$H_{t}\succeq\frac{1}{3e^{2}}M_{t}.$
The proof is deferred to Appendix C.2.1.
The following lemma is the elliptical potential lemma adapted to our setting.
Lemma C.3 (Elliptical potential for expected mean-centered uncertainty).
Let $M_{t}$ be defined as in Equation (C.2). Let $P_{t}(a\mid S_{t}^{(j)};\bm{\theta})$ denote the MNL probability for the subset of remaining actions in $S_{t}$ after removing the first j1j-1 actions according to the ranking $\sigma$. Then, for any $\lambda>0$, we have
$\sum_{t\notin{\mathcal{T}}^{w}}\min$ $\left\{1,\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S_{t};\bm{\theta}^{\star})\\ a\sim P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a\sim P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{M_{t}^{-1}}^{2}\right]\right\}$
$\leq 2d\log\left(1+\frac{KT}{d\lambda}\right).$
The proof is deferred to Appendix C.2.2.
The size of the set ${\mathcal{T}}^{w}$ is bounded as described in the following lemma.
Lemma C.4.
Let ${\mathcal{T}}^{w}=\big\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq\frac{1}{3\sqrt{2}K\beta_{T+1}(\delta)}\big\}$. Define $\kappa:=e^{-6B}$. Set $\lambda\geq 1$. Then, the size of the set ${\mathcal{T}}^{w}$ is bounded as follows:
$\left|{\mathcal{T}}^{w}\right|\leq\frac{288K^{4}}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right).$
The proof is deferred to Appendix C.2.3.
@@
$\leq\frac{1}{T}\sum_{t=1}^{T}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)+\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right),$ (C.3)
@@
To complete the proof, it remains to bound the first term in Equation (C.3).
@@
$=\frac{1}{T}\sum_{t\in{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\quad+\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{4B}{T}|{\mathcal{T}}^{w}|+\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$ (Assumption 1)
@@
$\leq\mathcal{O}\left(\frac{BK^{4}}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{T}{d\lambda}\right)\right)$ (Lemma C.4)
@@
$\quad+\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right).$ (C.4)
To further bound the last term of Equation (C.4), we get
@@
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}$ $\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{T+1}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ (Hölder’s ineq.)
@@
$\leq\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}\left\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right\|_{H_{T+1}}$ ($H_{T+1}\succeq H_{t}$)
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}.$ (Corollary C.1, with prob. $1-\delta$)
Define the sigmoid function $\mu(w)=\frac{1}{1+e^{-w}}$. Then, we have
$\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
@@
$=\sum_{t\notin{\mathcal{T}}^{w}}\frac{\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\right)}{\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\right)}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$=(1+e)^{2}\!\!\sum_{t\notin{\mathcal{T}}^{w}}\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\right)\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}},$
where the last equality holds due to the fact that
$\frac{1}{\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\right)}$
$\quad\quad=\frac{\Big(1+e^{\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}\Big)^{2}}{e^{\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}}$
$\quad\quad\leq\Big(1+e^{\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}}\Big)^{2}$ ($\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\geq 0$)
$\quad\quad\leq\left(1+e^{\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1})}\right)^{2}$ ($\widehat{\pi}_{T}(x_{t})=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widehat{\bm{\theta}}_{T+1}$)
$\quad\quad\leq\left(1+e^{\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\|_{H_{t}^{-1}}\|\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\|_{H_{T+1}}}\right)^{2}$ (Hölder’s ineq. and $H_{T+1}\succeq H_{t}$)